OpenStudy (anonymous):

The volume of a square pyramid whose base side is 9 cm and height is 5 cm is _____ cm³.

OpenStudy (anonymous):

405 cm^3

OpenStudy (anonymous):

Volume = 1/3 * BaseArea * Height So, Volume = 1/3 * 9 * 9 * 5 = 405/3

OpenStudy (anonymous):

@kiss69fist23 sorry, i misread the question.... andreh's solution is the correct one............ 1 Volume = --- * BaseArea * Height 3 1 405 Volume = --- * 9 * 9 * 5 = ----- = 135 cm^3 3 3